# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&\lVert m^{-1}(m_{T(x)}-m)P(f_{k})\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\frac{1}{\kappa}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$
$\leq C_{\zeta}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&C\lVert m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\lVert f_{k}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\frac{1}{\kappa}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$. (4.51)
@@
$\lVert\Gamma(f,f)\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\leq C\lVert f\rVert^{2}_{L^{\infty}_{a,0}(\Omega\times\mathbb{R}^{3})}.$ (4.52)
@@
$\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\leq C_{\alpha}\lVert T(x)-1\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}.$ (4.53)
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}(\mathcal{A}+\mathcal{B})$ (4.54)
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C_{\zeta}(\mathcal{A}+\mathcal{B})$
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (4.55)
$\leq C_{\zeta}(\mathcal{A}+\mathcal{B})+C_{\zeta}C\mathcal{B}(\mathcal{A}+\mathcal{B})+4CC^{2}_{\zeta}\frac{1}{\kappa}(\mathcal{A}+\mathcal{B})^{2}$ (4.56)
$\leq\Big(C_{\zeta}+C_{\zeta}C\mathcal{B}+4CC^{2}_{\zeta}\max\{1,\frac{1}{\kappa}\}(\mathcal{A}+\mathcal{B})\Big)(\mathcal{A}+\mathcal{B}).$
Choose $\delta<\min\{\frac{1}{4CC_{\zeta}},\frac{\min\{1,\kappa\}}{32CC^{2}_{\zeta}}\}$. Then, the proof is complete. Now, consider $f_{k+1}-f_{k}$ for $k\in\mathbb{N}$. Following a similar procedure, we obtain
@@
This implies that $\{f_{k}\}$ is a Cauchy sequence in $L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$. Hence, we obtain the solution by taking the limit $f=\displaystyle\lim_{k\to\infty}f_{k}$. Uniqueness follows by same argument. Further the upper bound of $f$, by (4.55), we have
$\lVert f\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C(\mathcal{A}+\mathcal{B})\leq C\delta.$ (4.57)
Chapter 5. Application to Thermal transpiration and thermomolecular pressure difference
@@
In this chapter, we prove the Theorem 1.2. Without loss of generality, we assume that $a=0$ and $b=L$. To simplify the calculation and to obtain a more explicit estimate, we prove Theorem 1.2 when $\Omega$ satisfies
@@
$\begin{aligned} F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{v^{2}}{2}},&&\,\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}\cap\{x_{3}<0\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{5/2}_{2}}e^{-\frac{v^{2}}{2T_{2}}},&&\,\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}\cap\{x_{3}>L\},\\ F(x,v)&=m_{T(x)}\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\,\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}$ (5.1)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=0,&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (5.3)
The proof appears in the [36]. Now, the equation (1.5) becomes
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=\frac{2}{\kappa}Q(m_{T_{q}},f)+\frac{1}{\kappa}Q(f,f)+\frac{1}{\kappa}Q(m_{T_{q}},m_{T_{q}}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,a},\\ f&=\frac{1}{\sqrt{2\pi}}\left(\frac{1}{\sqrt{T_{2}}}-1\right)\frac{1}{2\pi T_{2}^{2}}e^{-\frac{|v|^{2}}{2T_{2}}}\coloneq-C_{T_{2}}m_{T_{2}},&&\quad\text{on}\quad\partial\Omega^{-}_{1,b},\\ f&=\frac{1}{(2\pi)T^{2}(x)}e^{-\frac{|v|^{2}}{2T(x)}}\int_{v\cdot n(x)>0}f(x,v)v\cdot n(x)dv,&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (5.4)
@@
$u(x)\coloneq\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv,\quad U(z)\coloneq\int_{\{x_{1}^{2}+x_{2}^{2}\leq 1,x_{3}=z\}}u(x)dS.$ (5.5)
@@
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\partial\Omega_{1,a}^{-}\right\},$ (5.6)
@@
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}<\cos\theta\leq 1\right\},$ (5.7)
@@
$\cos^{-1}\left(\frac{z_{0}-L}{\sqrt{(z_{0}-L)^{2}+4\sin^{2}\phi}}\right)=\pi\quad\text{and}\quad\cos^{-1}\left(\frac{z_{0}}{\sqrt{z_{0}^{2}+4\sin^{2}\phi}}\right)=0$ (5.8)
@@
Note that the $f$ in 5.4 has the form
$f(x,v)$ $=-\textbf{1}_{\partial\Omega^{-}_{1,b}}(q(x,v))C_{T_{2}}m_{T_{2}}$ (5.9)
@@
We now use a procedure similar to the one in Chapter 4 . When tracing back $n+1$ time for $n\geq 1$, we obtain
$\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv$ $=\int_{\mathbb{R}^{3}}v_{3}f(x,v)dv$ (5.10)
@@
Now, we estimate each term in (5.10). The definition of $A_{1}$ to $D$ appearing in the last line of (5.10) are:
$A_{1}$ $=\int_{\mathbb{R}^{3}}v_{3}\textbf{1}_{\partial\Omega^{-}_{1,b}}gdv,$ (5.11)
$A_{2}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g)dv,$ (5.12)
$B_{j}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g),$ (5.13)
$C_{1}$ $=\int_{\mathbb{R}^{3}}v_{3}h[f]dv,$ (5.14)
$C_{2}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(h[f])dv+\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(h[f])dv,$ (5.15)
$D$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{n+1}\circ\overline{P}(f)dv.$ (5.16)
For $A_{1}$, using (5.7). we obtain
@@
$\Theta(x)$ $=2\int_{-\pi/2}^{\pi/2}\int^{\pi}_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+4\sin^{2}\phi}}\right)}\cos\phi\sin^{2}\theta d\theta d\phi,$ (5.17)
$\overline{\Theta}_{1}(x)$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\Theta dv$ (5.18)
@@
$\overline{\Theta}_{2}(x)$ $=\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}(\Theta(x)).$ (5.19)
@@
$U(0)=\int_{\{x_{1}^{2}+x_{2}^{2}\leq 1,x_{3}=0\}}u(x)dS\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (5.20)
@@
For the general case, the terms $A_{1},A_{2}$ and $B_{j}$ are very difficult to write down in explicit form, since (5.7) depends on $\partial\Omega\cap\{x_{3}=0\}$ and $\partial\Omega\cap\{x_{3}=L\}$. However, we can still prove that $\Omega$ is a completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$, and the proof of the general case proceed in the same way as before.
@@
Since $\Omega$ is an open set, there exist $\rho_{1},\rho_{2}>0$ and $a_{1},a_{2}\in\mathbb{R}^{3}$ such that
$B(a_{1},2\rho_{1})\cap\{x_{3}=0\}\subset\Omega\cap\{x_{3}=0\}\quad\text{and}\quad B(a_{2},2\rho_{2})\cap\{x_{3}=L\}\subset\Omega\cap\{x_{3}=L\}.$ (5.21)
Without loss of generality, we may assume that $a_{1}=0$ and check the definition 4.1 when $x\in\partial\Omega_{2}$ such that $x=(x_{1},0,x_{3})$ with $x_{1}>\rho_{1}$. We classify $v\in\mathbb{R}^{3}$ as follows
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\left[B(a_{1},2\rho_{1})\cap\{x_{3}=0\}\right]^{-}\right\},$
$A_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\left[B(a_{2},2\rho_{2})\cap\{x_{3}=0\}\right]^{-}\right\},$

### v2

@@ 4.1. The existence of solution to (80)
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ $\leq C_{\zeta}\max\{1,\kappa^{-1}\}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&\lVert m^{-1}(m_{T(x)}-m)P(f_{k})\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\kappa^{-1}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\begin{pmatrix}[l]&\lVert f_{in}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{1})}\\ +&C\lVert m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\lVert f_{k}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\\ +&\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\\ +&\max\{1,\kappa^{-1}\}\lVert\Gamma(f_{k},f_{k})\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\end{pmatrix}$. (128)
@@
$\lVert\Gamma(f,f)\rVert_{L^{\infty}_{a,-1}(\Omega\times\mathbb{R}^{3})}\leq C\lVert f\rVert^{2}_{L^{\infty}_{a,0}(\Omega\times\mathbb{R}^{3})}.$ (129)
@@
$\lVert\frac{1}{\sqrt{2\pi}}m^{-1/2}(m_{T(x)}-m)\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{-}_{2})}\leq C_{\alpha}\lVert T(x)-1\rVert_{L^{\infty}(\partial\Omega^{-}_{2})}.$ (130)
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq 2C_{\zeta}\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B})$ (131)
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C_{\zeta}\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B})$
@@
$\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f_{k+1}\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}$ (132)
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\Big[(\mathcal{A}+\mathcal{B})+C\mathcal{B}(\mathcal{A}+\mathcal{B})+4CC_{\zeta}\max\{1,\kappa^{-3}\}(\mathcal{A}+\mathcal{B})^{2}\Big]$ (133)
$\leq C_{\zeta}\max\{1,\kappa^{-1}\}\Big[1+C\mathcal{B}+4CC_{\zeta}\max\{1,\kappa^{-2}\}(\mathcal{A}+\mathcal{B})\Big](\mathcal{A}+\mathcal{B}).$
We choose
$\delta<\min\left\{\frac{1}{4C},\frac{1}{16CC_{\zeta}\max\{1,\kappa^{-2}\}}\right\}.$
Then, the proof is complete. Now, consider $f_{k+1}-f_{k}$ for $k\in\mathbb{N}$. Following a similar procedure, we obtain
@@
This implies that $\{f_{k}\}$ is a Cauchy sequence in $L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})$. Hence, we obtain the solution by taking the limit $f=\lim_{k\to\infty}f_{k}$. Uniqueness follows by same argument. Further the upper bound of $f$, by (132), we have
$\lVert f\rVert_{L^{\infty}_{\alpha,0}(\Omega\times\mathbb{R}^{3})}+\lVert f\rVert_{L^{\infty}_{\alpha,0}(\partial\Omega^{+}_{2})}\leq C\max\{1,\kappa^{-1}\}(\mathcal{A}+\mathcal{B}).$ (134)
Section 5. Application to Thermal transpiration and thermomolecular pressure difference
@@
In this section, we prove the Theorem 1.2. Without loss of generality, we assume that $a=0$ and $b=L$. To simplify the calculation and to obtain a more explicit estimate, we prove Theorem 1.2 when $\Omega$ satisfies
@@
$\begin{aligned} F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{v^{2}}{2}},&&\,\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}\cap\{x_{3}<0\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{5/2}_{2}}e^{-\frac{v^{2}}{2T_{2}}},&&\,\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}\cap\{x_{3}>L\},\\ F(x,v)&=m_{T(x)}\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\,\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}$ (135)
@@
$\left\{\begin{aligned} v\cdot\nabla_{x}F&=0,&&\text{in}\quad\Omega\times\mathbb{R}^{3},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}}e^{-\frac{|v|^{2}}{2}},&&\text{on}\quad\partial\Omega^{-}_{1,a}=\partial\Omega^{-}_{1}\cap\{x_{3}<a\},\\ F(x,v)&=\frac{1}{(2\pi)^{3/2}T^{2}_{2}}e^{-\frac{|v|^{2}}{2T_{2}}},&&\text{on}\quad\partial\Omega^{-}_{1,b}=\partial\Omega^{-}_{1}\cap\{x_{3}>b\},\\ F(x,v)&=m_{T}(x,v)\int_{v\cdot n(x)>0}F(x,v)v\cdot n(x)dv,&&\text{on}\quad\partial\Omega^{-}_{2}=\partial\Omega^{-}\setminus\left(\partial\Omega^{-}_{1,a}\cup\partial\Omega^{-}_{1,b}\right).\end{aligned}\right$. (137)
The proof appears in the [36]. Now, the equation (7) becomes
$\left\{\begin{aligned} v\cdot\nabla_{x}f&=\frac{2}{\kappa}Q(m_{T_{q}},f)+\frac{1}{\kappa}Q(f,f)+\frac{1}{\kappa}Q(m_{T_{q}},m_{T_{q}}),&&\quad\text{in}\quad\Omega\times\mathbb{R}^{3},\\ f&=0,&&\quad\text{on}\quad\partial\Omega^{-}_{1,a},\\ f&=\frac{1}{\sqrt{2\pi}}\left(\frac{1}{\sqrt{T_{2}}}-1\right)\frac{1}{2\pi T_{2}^{2}}e^{-\frac{|v|^{2}}{2T_{2}}}\coloneq-C_{T_{2}}m_{T_{2}},&&\quad\text{on}\quad\partial\Omega^{-}_{1,b},\\ f&=\frac{1}{(2\pi)T^{2}(x)}e^{-\frac{|v|^{2}}{2T(x)}}\int_{v\cdot n(x)>0}f(x,v)v\cdot n(x)dv,&&\quad\text{on}\quad\partial\Omega^{-}_{2}.\end{aligned}\right$. (138)
@@
$u(x)\coloneq\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv,\quad U(z)\coloneq\int_{\{x_{1}^{2}+x_{2}^{2}\leq 1,x_{3}=z\}}u(x)dS.$ (139)
@@
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,(q(x,v),v)\in\partial\Omega_{1,a}^{-}\right\},$ (140)
@@
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,\frac{x_{3}}{\sqrt{x_{3}^{2}+r^{2}}}<\cos\theta\leq 1\right\},$ (141)
@@
$\cos^{-1}\left(\frac{z_{0}-L}{\sqrt{(z_{0}-L)^{2}+4\sin^{2}\phi}}\right)=\pi\quad\text{and}\quad\cos^{-1}\left(\frac{z_{0}}{\sqrt{z_{0}^{2}+4\sin^{2}\phi}}\right)=0$ (142)
@@
Note that the $f$ in 138 has the form
$f(x,v)$ $=-\textbf{1}_{\partial\Omega^{-}_{1,b}}(q(x,v))C_{T_{2}}m_{T_{2}}$ (143)
@@
We now use a procedure similar to the one in section 4 . When tracing back $n+1$ time for $n\geq 1$, we obtain
$\int_{\mathbb{R}^{3}}v_{3}F(x,v)dv$ $=\int_{\mathbb{R}^{3}}v_{3}f(x,v)dv$ (144)
@@
Now, we estimate each term in (144). The definition of $A_{1}$ to $D$ appearing in the last line of (144) are:
$A_{1}$ $=\int_{\mathbb{R}^{3}}v_{3}\textbf{1}_{\partial\Omega^{-}_{1,b}}gdv,$ (145)
$A_{2}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g)dv,$ (146)
$B_{j}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(\textbf{1}_{\partial\Omega^{-}_{1,b}}g),$ (147)
$C_{1}$ $=\int_{\mathbb{R}^{3}}v_{3}h[f]dv,$ (148)
$C_{2}$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\overline{P}(h[f])dv+\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}\circ\overline{P}(h[f])dv,$ (149)
$D$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{n+1}\circ\overline{P}(f)dv.$ (150)
For $A_{1}$, using (141). we obtain
@@
$\Theta(x)$ $=2\int_{-\pi/2}^{\pi/2}\int^{\pi}_{\cos^{-1}\left(\frac{x_{3}-L}{\sqrt{(x_{3}-L)^{2}+4\sin^{2}\phi}}\right)}\cos\phi\sin^{2}\theta d\theta d\phi,$ (151)
$\overline{\Theta}_{1}(x)$ $=\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\Theta dv$ (152)
@@
$\overline{\Theta}_{2}(x)$ $=\sum_{j=1}^{n}\int_{\mathbb{R}^{3}}v_{3}J^{T}_{\partial\Omega^{-}_{2}}\hat{P}^{j}(\Theta(x)).$ (153)
@@
$U(0)=\int_{\{x_{1}^{2}+x_{2}^{2}\leq 1,x_{3}=0\}}u(x)dS\geq C\left(1-\frac{1}{\sqrt{T_{2}}}\right).$ (154)
@@
For the general case, the terms $A_{1},A_{2}$ and $B_{j}$ are very difficult to write down in explicit form, since (141) depends on $\partial\Omega\cap\{x_{3}=0\}$ and $\partial\Omega\cap\{x_{3}=L\}$. However, we can still prove that $\Omega$ is a completely unsealed set with respect to $\partial\Omega_{1,a}\cup\partial\Omega_{1,b}$, and the proof of the general case proceed in the same way as before.
@@
Since $\Omega$ is an open set, there exist $\rho>0$ and $a_{1}=(a_{1,1},a_{1,2},0),a_{2}=(a_{2,1},a_{2,2},L)\in\mathbb{R}^{3}$ such that
$B(a_{1},\rho)\cap\{x_{3}=0\}\subset\Omega\cap\{x_{3}=0\}\quad\text{and}\quad B(a_{2},\rho)\cap\{x_{3}=L\}\subset\Omega\cap\{x_{3}=L\}.$ (155)
Then we define
$\Pi=B(a_{1},\rho)\cap\{x_{3}=0\}\cup B(a_{2},\rho)\cap\{x_{3}=L\}.$ (156)
We show that $\Omega\cap\{0\leq x_{3}\leq L\}$ is completely unsealed with respect to $\Pi$. We can then deduce the lemma from this conclusion. For any fixed $x\in\Xi\coloneq\partial(\Omega\cap\{0\leq x_{3}\leq L\})\setminus(\{x_{3}=0\}\cup\{x_{3}=L\})$, without loss of generality, we may assume that $a_{1}=(0,0,0)$ and $x=(x_{1},0,x_{3})$. Then we classify $v\in\mathbb{R}^{3}$ as follows.
$A_{1}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,q(x,v)\in B(0,\rho)\cap\{x_{3}=0\}\right\},$
$A_{2}$ $=\left\{(v_{1},v_{2},v_{3})\in\mathbb{R}^{3}\,\Big|\,q(x,v)\in B(a_{2},\rho)\cap\{x_{3}=L\}\right\},$
$A_{3}$ $=\mathbb{R}^{3}\setminus(A_{1}\cup A_{2}).$
Let $v=(\rho\sin\theta\cos\phi,\rho\sin\theta\sin\phi,\rho\cos\theta)$. Then the sets $A_{1}$ and $A_{2}$ can be written more explicitly as follow.
$A_{1}=$ $\left\{(\rho,\phi,\theta)\in\mathbb{R}_{+}\times[0,2\pi]\times[0,\frac{\pi}{2}]\,\Big|0\leq\alpha_{1}^{2}+\gamma_{1},\quad\frac{\alpha_{1}-\sqrt{\alpha_{1}^{2}+\gamma_{1}}}{x_{3}}\leq\tan\theta\leq\frac{\alpha_{1}+\sqrt{\alpha_{1}^{2}+\gamma_{1}}}{x_{3}}\right\}$ (157)
$A_{2}=$ $\left\{(\rho,\phi,\theta)\in\mathbb{R}_{+}\times[0,2\pi]\times[\frac{\pi}{2},\pi]\,\Big|0\leq\alpha_{2}^{2}+\gamma,\quad\frac{-\alpha_{2}-\sqrt{\alpha_{2}^{2}+\gamma_{2}}}{L-x_{3}}\leq\tan\theta\leq\frac{-\alpha_{2}+\sqrt{\alpha_{2}^{2}+\gamma_{2}}}{L-x_{3}}\right\}$ (158)
where
$\begin{aligned} &\alpha_{1}=x_{1}\cos\phi,\,&&\gamma_{1}=\rho^{2}-x_{1}^{2}\\ &\alpha_{2}=[x_{1}-a_{2,1}]\cos\phi-a_{2,2}\sin\phi,\,&&\gamma_{2}=\rho^{2}-[(x_{1}-a_{2,1})^{2}+a_{2,2}^{2}].\end{aligned}$
Let
$W_{1}(x)\coloneq\int_{u\cdot n(x)>0}1_{\partial(\Omega\cap\{0\leq x_{3}\leq L\})\setminus A_{1}}(q(x,u))m_{T(q(x,u))}u\cdot n(x)du=1-\int_{u\cdot n(x)>0}1_{A_{1}}(u)m_{T(q(x,u))}u\cdot n(x)du$